The solution to the system of equations
x + 3y = 12
4x - y = -17 is?

Answer by scott8148(6628)

x=12-3y ... 4(12-3y)-y=-17 ... 48-12y-y=-17 ... -13y=-65 ... y=5 ... x=12-3(5) ... x=-3